### v1

# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

Abstract
LiDAR segmentation is crucial for autonomous driving systems. The recent range-view approaches are promising for real-time processing. However, they suffer inevitably from corrupted contextual information and rely heavily on post-processing techniques for prediction refinement. In this work, we propose a simple yet powerful FRNet that restores the contextual information of the range image pixels with corresponding frustum LiDAR points. Firstly, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is crucial for point-level predictions. Next, a frustum-point fusion module is introduced to update per-point features hierarchically, which enables each point to extract more surrounding information via the frustum features. Finally, a head fusion module is used to fuse features at different levels for final semantic prediction. Extensive experiments on four popular LiDAR segmentation benchmarks under various task setups demonstrate our superiority. FRNet achieves competitive performance while maintaining high efficiency. The code is publicly available.
1 Introduction
Figure 1: Scalability comparisons among state-of-the-art LiDAR segmentation models on the SemanticKITTI [5] leaderboard. The size of the circular represents the number of model parameters. Our FRNet achieves competitive performance with current arts while maintaining satisfactory efficiency for real-time processing.
Real-time and accurate LiDAR semantic segmentation is crucial in autonomous driving, robotics, and other safety-critical applications [18]. Many recent approaches have made significant progress in this area, which represent LiDAR point clouds from different perspectives, including point view [48, 21], voxel view [47, 65], pseudo-image view [50, 35, 61], and multi-view representation [31, 54, 2], depending on different pre-processing techniques.
Point-view methods manipulate the origin point cloud to preserve raw spatial information, following the paradigm of PointNet++ [40]. Nevertheless, they require computationally expensive neighborhood searches to construct local neighborhoods, and the computation overhead is considerable for large scene point clouds. Voxel-view methods transform scatter points into regular voxel representations and apply 3D convolutions for label prediction, but the computation grows cubically when voxel resolution increases. Multi-view methods extract features from different representations separately and fuse these features from different levels for more accurate prediction. Consequently, the computation of all representations is also gathered. Albeit these representation methods achieve some satisfactory performance, their computation overhead hinders them from real-time online applications.
Recently, pseudo-image methods propose a simple yet efficient intermediate representation for LiDAR processing, such as bird’s-eye-view [61, 1] and range-view [50, 51, 43, 4, 24]. Under such representations, some popular image segmentation approaches [33, 62] can be directly applied to such pseudo-images to balance the efficiency and accuracy of segmentation. However, the 3D-to-2D projection procedure tends to result in corrupted neighborhood contextual information, decreasing the precision. Detailed analysis of this problem can be summarized into two aspects.
Firstly, some 3D points are discarded in the projection procedure. Under such a paradigm, only the projected points are involved in convolutions, while the contextual information of discarded points is absent. Consequently, the pixel in the pseudo-image can hardly preserve the high-fidelity 3D geometric context. More importantly, these discarded 3D points still need to predict 3D labels but no features correspond to them. To reconstruct these labels, some post-processing techniques, such as CRF [22] and KNN [35], are used for these points based on 2D predictions. However, these techniques are unsupervised and fail to extract the local features of discarded points in a learnable manner, which significantly limits the improvement of pseudo-image methods.
Secondly, the pseudo-image suffers from many empty pixels due to the sparsity of LiDAR points. Under such a non-compact representation, the projected points will extract much noisy information as the convolutions go deeper. In summary, pseudo-image methods fail to preserve the contextual information of all the points, which is significant for LiDAR segmentation. So, it is important for points and pseudo-images to cooperate with each other. Although existing multi-view methods have successfully preserved the context of all points via fusing features of different representations, they are always computationally expensive.
To this end, we study a scalable Frustum-Range Network (FRNet) for LiDAR segmentation, which can preserve the contextual information of all points through the cooperation of point and pseudo-image and still maintain high efficiency. Specifically, FRNet comprises three main components. To preserve all points and predict semantic labels for these points in an end-to-end manner, a Frustum Feature Encoder is used to group all points with the same frustum region into the corresponding range-view pixel in relation to the range image via several multi-layer perceptions. The point features can be pooled to represent the frustum region and formatted into a 2D representation, followed by traditional convolutions. Then, a Frustum-Point (FP) Fusion Module is introduced to update the features of each point hierarchically during each stage of the convolutions, which consists of frustum-to-point fusion to update per-point features and point-to-frustum fusion to enhance frustum features. In this way, all the points extract larger local features based on the frustum region. Finally, a Fusion Head module is designed to utilize different level features to generate individual features for each point, so that the results can be predicted end-to-end without any post-processing techniques. As shown in Fig. 1, the proposed FRNet achieves great improvements among range-view methods while still maintaining high efficiency.
To alleviate empty pixels in 2D representation, we introduce a new range interpolation augmentation method that adds several potential points based on the information of surrounding points. Specifically, we first transform the point clouds into range images, as done in previous works. For empty pixels in the range images, we then calculate their potential coordinates when both the nearest left and right pixels have projected points as they often share similar attributes. With this augmentation, the 2D representation will have fewer empty pixels, allowing convolutions to learn more semantically coherent information.
The key contributions are summarized as follows:
  • We introduce a novel Frustum-Range representation for LiDAR segmentation. In this representation, each pixel in the 2D representation is treated as a corresponding frustum region in the point clouds. All points within the same frustum region are preserved for semantic prediction without the need for any post-processing techniques.
  • We propose a novel range interpolation augmentation that synthesizes possibly lost LiDAR points based on surrounding information, helping to fill empty pixels and learn more coherent range-view representations.
  • Extensive experiments across four LiDAR segmentation benchmarks demonstrate that FRNet can achieve the optimal trade-off between efficiency and accuracy.
2 Related Work
Scalable LiDAR Segmentation. Although Point-view [21, 48], voxel-view [47, 65], and multi-view [31, 32] have achieved great success in LiDAR segmentation, they are often time-consuming and cannot achieve real-time processing. Recently, various methods proposed to project LiDAR points into range images along inclination and azimuth directions to balance the efficiency and accuracy of segmentation. SqueezeSeg [50] and SqueezeSegV2 [51] used the lightweight model SqueezeNet [22] to retain information. SqueezeSegV3 [53] introduced a Spatially-Adaptive Convolution module to adjust kernel weights based on the locations of range images. RangeNet++ [35] integrated Darknet into SqueezeSeg [50] and proposed an efficient KNN post-processing for segmentation. RangeViT [4] and RangeFormer [24] introduced popular transformer blocks to extract local and global information of the range images. These methods only learn the features of projected points while ignoring the occluded points.
LiDAR Post-Processing. Prior works have proposed various post-processing techniques to reconstruct the semantic labels of occluded points based on 2D results. Many of these methods follow the CRF [50] or KNN [35] approach to smooth the semantic results of occluded points. FIDNet [64] proposed a Nearest Label Assignment (NLA) method for label smoothing. Both of these techniques are unsupervised and can not extract the features of all points in a learnable manner. RangeFormer [24] created sub-clouds from full point clouds and inferred each subset in a supervised manner, but it still relies on 2D prediction results. In contrast, our method involves the frustum region and predicts the semantic label of each point directly without any information loss, making it a fully end-to-end approach.
LiDAR Data Augmentation. Recent works have explored different data augmentation techniques on point clouds [7, 60], which can help improve performance when it comes to indoor scene understanding but has limited generalization to outdoor scenes. Mix3D [36] created new training samples by combining two scenes out of context. PolarMix [52] proposed mixing two LiDAR scenes based on yaw angles and cropping instances from one scene to another, while LaserMix [26] proposed exchanging labeled scans with unlabeled scans along the pitch axis for efficient semi-supervised learning. RangeFormer [24] presented a lightweight augmentation method that only operated on projected points in range images, disregarding occluded points. In this work, we prevent a novel augmentation technique based on our proposed frustum region, which preserves all points within the frustum region, encompassing techniques such as FrustumMix and Range-Interpolation.
Figure 2: A pilot study on the performance degradation of post-processing in existing range-view methods [35, 12, 64, 8] on SemanticKITTI [5]. We choose various $K$ values as hyperparameters in KNN post-processing. Compared to their performance at 2D, a severe drop in performance occurs with different $K$ values.
3 FRNet: A Scalable LiDAR Segmentor
3.1 Pilot Study
Figure 3: Architecture overview. The proposed FRNet comprises three main components: 1) Frustum Feature Encoder is used to embed per-point features in the frustum region. 2) Frustum-Point (FP) Fusion Module is used to update per-point features hierarchically at each stage of the 2D backbone. 3) Fusion Head fuses different levels of features to predict final results.
LiDAR segmentation aims to assign semantic labels to each point in the point clouds. A common approach is to perform operations directly on the LiDAR points without any pre-processing, following the paradigm of PointNet++ [40]. However, point-based operations often require extensive computations and can yield sub-optimal performance due to the sparsity of outdoor scenes. To address this issue, many recent works propose projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency. Given point clouds with coordinates $(x,y,z)$ and the expected projected resolution $(H,W)$ of range images, each point can be mapped to its correspondence coordinates $(u,v)$ in the range image via the following transformation:
$\binom{u}{v}=\binom{\frac{1}{2}[1-\arctan(y,x)\pi^{-1}]W}{[1-(\arcsin(zd^{-1})+f_{up})f^{-1}]H}~{},$ (1)
where $d=\sqrt{x^{2}+y^{2}+z^{2}}$ represents the depth of a point, $f=|f_{up}|+|f_{down}|$ represents the vertical field-of-view of the sensor, and $f_{up}$ and $f_{down}$ are the pitch angles at the upward and downward directions, respectively. However, such a projection process can cause the loss of occluded points and often requires post-processing techniques to reduce the blurry prediction of lost points [24].
We conduct a pilot study on different range-view methods, including RangeNet++ [35], SalsaNeXt [12], FIDNet [64], and CENet [8], and compare their performance with different KNN post-processing parameters and 2D performance. As shown in Fig. 2, the post-processing procedures are often unsupervised, and the hyperparameters can significantly influence the final performance. To address this issue, we propose a scalable Frustum-Range Network (FRNet) that can directly predict semantic labels for each point while maintaining high efficiency.
3.2 Frustum-Range Architecture
Given a point cloud $\mathbf{P}$ acquired by the LiDAR sensor, FRNet is expected to learn a feed-forward network $\Phi$ to predict semantic labels $\mathbf{Y}$ for each point:
$\mathbf{Y}=\Phi(\mathbf{P},\theta)~{},$ (2)
where $\theta$ denotes the learnable parameters in the network. As shown in Fig. 3, FRNet comprises a Frustum Feature Encoder that carries out per-point feature extraction, a Frustum-Point Fusion Module that fuses per-point features and frustum features hierarchically, and a Fusion Head Module that predicts semantic labels for each point.
Frustum Feature Encoder. LiDAR segmentation involves predicting semantic labels for each point, so it is important to extract individual features for each point. Before converting point clouds into 2D representations, Frustum Feature Encoder first extracts per-point features via several MLPs, which is crucial for accurate label prediction. For points $\mathbf{P}_{i}=\{p_{i}^{1},p_{i}^{2},...,p_{i}^{N}\}$ in the $i$-th frustum region, the points are manually embedded with $(p_{i}^{k},p_{i}^{k}-p_{i}^{mean})$, where $p_{i}^{k}$ is the $k$-th point in the frustum, $N$ is the total number of points in the frustum region, and $p_{i}^{mean}=\frac{1}{N}\sum_{k=1}^{N}p_{i}^{k}$ represents the mean information of the frustum region. The per-point features are then obtained as follows:
$f_{i}^{k}=\text{MLP}([p_{i}^{k};p_{i}^{mean}])~{},$ (3)
[…]
Frustum-Point Fusion Module. As shown in Fig. 4, the Frustum-Point (FP) Fusion Module comprises two components: 1) A Frustum-to-Point fusion that hierarchically updates per-point features, and 2) A Point-to-Frustum fusion to fuse individual features of each point within its corresponding frustum region. Specifically, for the point features $F_{p}$ and frustum features $F_{f}$, $F_{f}$ is back-project to the point-level features $F_{p}^{f}$ based on the projection index. This means that points within the same frustum region share the same frustum features. Then, the two features are concatenated together, followed by an MLP to update per-point features:
$\hat{F}_{p}=\text{MLP}([F_{p}^{f};F_{p}])~{}.$ (4)
Next, the updated per-point features $\hat{F}_{p}$ are pooled into the corresponding frustum region to further enhance the frustum features $F_{f}$ in a residual-attentive manner. The pooled features are denoted as $F_{f}^{p}$, which are calculated with the max operation based on the frustum region. The pooled features $F_{f}^{p}$ are first concatenated with $F_{f}$, followed by a simple convolutional layer to reduce the number of channels:
$F^{fuse}=g([F_{f}^{p};F_{f}],\theta_{1})~{},$ (5)
where $g(\cdot)$ is the mapping function and $\theta_{1}$ is the learnable parameters. Finally, the fused features are used to further enhance the frustum features in a residual-attentive manner:
$\hat{F}_{f}=F_{f}+\sigma(h(F^{fuse},\theta_{2}))\odot F^{fuse}~{},$ (6)
where $\sigma$ represents the sigmoid function, $h$ is the mapping function, $\theta_{2}$ is the learnable parameters, and $\odot$ indicates element-wise multiplication. The output features $\hat{F}_{f}$ are then used to extract larger local frustum features with subsequent convolutions.
Fusion Head Module. Since the aim of FRNet is to classify categories for each point, the most straightforward method is to directly use the updated per-point features after the FP Fusion Module for prediction. Although this can also achieve great performance, we observe that this is not the optimal solution. The output of our FP Fusion Module consists of two features: per-point features and frustum features represented in a 2D manner, and our prediction head expects to fully utilize both of these two features. Additionally, the per-point features generated from the Frustum Feature Encoder Module directly embed coordinate information, which can also provide more low-level geometric information. The Fusion Head Module progressively fuses three different level features. Firstly, the frustum features $F_{f}$ are mapped to point-level features $F_{f}^{p}$. It is worth noting that points in the same frustum regions share the same features at this stage. Then, we map the features $F_{f}^{p}$ with MLPs to ensure they share the same channels as point features. Finally, the add operation is applied to fuse the two features. The same operation is also used to fuse with low-level point features $F_{p}^{l}$. The overall operation can be formulated as:
$F_{out}=\text{MLP}(\text{MLP}(F_{f}^{p})+F_{p})+F_{p}^{l}~{}.$ (7)
Here, $F_{out}$ is used to generate the final semantic scores for each point.
Optimization. To improve the optimization of FRNet, we not only utilize the commonly used point-level supervised loss function but also supervise the frustum-level features with generated pseudo frustum labels. Typically, a frustum region may contain multiple points, and these points often do not belong to the same category. Based on this observation, we generate a pseudo label for each frustum in a statistical manner. We count the frequency of each category in the frustum region and choose the category with the highest frequency as the pseudo label for the frustum. Finally, the frustum-level features will be supervised using the generated pseudo labels. The overall loss can be calculated as:
$L=L_{p}+\lambda L_{f}~{},$ (8)
where $L_{p}$ and $L_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision. $L_{p}$ is calculated by cross-entropy loss and $L_{f}$ consists of cross-entropy loss, Lovasz-Softmax loss [6], and boundary loss [42].
Figure 4: Frustum-Point fusion module comprises two steps: 1) A Frustum-to-Point fusion to update per-point features. 2) A Point-to-Frustum fusion to update frustum features.
3.3 Frustum-Range Operators
Data augmentation enables the model to learn more robust representations. Commonly used techniques in LiDAR segmentation often operate on the global view, including rotation, flipping, scaling, jittering, etc. In this work, we introduce two novel augmentations based on our proposed frustum-range view representation.
FrustumMix. Different from the previous LaserMix [26] and PolarMix [52] augmentations which mix two LiDAR scenes in the point level, we propose a FrustumMix to mix two scenes based on frustum units. Points in the scan will be first grouped into different frustum regions along inclination and azimuth directions. We then randomly split frustum regions equally from two directions and switch the corresponding frustum regions from two scans.
[…]
Method Venue Representation Param SemanticKITTI nuScenes ScribbleKITTI SemanticPOSS
FPS normal-$\uparrow$ Val normal-$\uparrow$ Test normal-$\uparrow$ Val normal-$\uparrow$ Test normal-$\uparrow$ mIoU normal-$\uparrow$ mAcc normal-$\uparrow$ mIoU normal-$\uparrow$ mAcc normal-$\uparrow$
MinkUNet [10] CVPR’19 Sparse Voxel $\bullet$ $21.7$ M
$\mathbf{9.1}$
$63.7$
$55.0$
$\mathbf{53.1}$
$\mathbf{68.3}$
SPVNAS [47] ECCV’20 Sparse Voxel $\bullet$ $21.8$ M $8.9$
$\underline{66.4}$
$\underline{56.9}$
$48.4$
$61.5$
Cylinder3D [65] CVPR’21 Sparse Voxel $\bullet$ $55.9$ M
$6.2$
$\mathbf{65.9}$
$\mathbf{67.8}$
$\mathbf{76.1}$
$\mathbf{77.9}$
$\mathbf{57.0}$
$\underline{52.9}$
$\underline{64.9}$
RPVNet [54] ICCV’21 Multi-View $\bullet$ $24.8$ M $7.9$
$65.5$
$70.3$
$77.6$
2DPASS [56] ECCV’22 Multi-View $\bullet$ $26.5$ M
$\mathbf{8.4}$
$69.3$
$72.2$
$\mathbf{79.4}$
$80.8$
SphereFormer [27] CVPR’23 Multi-View $\bullet$ $32.3$ M
$4.9$
$67.8$
$74.8$ $78.4$
$\underline{81.9}$
UniSeg [32] ICCV’23 Multi-View $\bullet$ $147.6$ M
$6.9$
$\mathbf{71.3}$
$\mathbf{75.2}$
$\mathbf{83.5}$
RangeNet++ [35] IROS’19 Range View $\bullet$ $50.4$ M
$12.0$
$50.3$
$52.2$
$65.5$
$44.6$
$57.8$
$30.9$
FIDNet [64] IROS’20 Range View $\bullet$ $6.1$ M
$31.8$
$58.9$
$59.5$
$54.1$
$65.4$
$46.4$
CENet [8] ICME’22 Range View $\bullet$ $6.8$ M $33.4$
$62.6$
$64.7$
$55.7$
$66.8$
$50.3$
RangeViT [4] CVPR’23 Range View $\bullet$ $23.7$ M
$10.0$
$60.7$
$64.0$
$75.2$
$53.6$
$66.5$
RangeFormer [24] ICCV’23 Range View $\bullet$ $24.3$ M
$6.2$
$67.6$
$\mathbf{73.3}$
$78.1$
$80.1$
$63.0$
Fast-FRNet Ours Frustum-Range $\bullet$ $7.5$ M
$\mathbf{33.8}$
$67.1$
$\underline{72.5}$
$78.8$ $82.1$
$62.4$
$71.2$ $52.5$ $67.1$
FRNet Ours Frustum-Range $\bullet$ $10.0$ M
$29.1$
$\mathbf{68.7}$
$\mathbf{73.3}$
$\mathbf{79.0}$
$\mathbf{82.5}$
$\mathbf{63.1}$
$\mathbf{72.3}$
$\mathbf{53.5}$
$\mathbf{68.1}$
[…]
Method Venue Backbone SemanticKITTI nuScenes ScribbleKITTI
$\mathbf{1\%}$ $\mathbf{10\%}$ $\mathbf{20\%}$ $\mathbf{50\%}$ $\mathbf{1\%}$ $\mathbf{10\%}$ $\mathbf{20\%}$ $\mathbf{50\%}$ $\mathbf{1\%}$ $\mathbf{10\%}$ $\mathbf{20\%}$ $\mathbf{50\%}$
Sup.-only - FIDNet [64] $36.2$ $52.2$ $55.9$ $57.2$ $38.3$ $57.5$ $62.7$ $67.6$ $33.1$ $47.7$ $49.9$ $52.5$
LaserMix [26] CVPR’23 $43.4$ $58.8$ $59.4$ $61.4$ $49.5$ $68.2$ $70.6$ $73.0$ $38.3$ $54.4$ $55.6$ $58.7$
Sup.-only - Cylinder3D [65] $45.4$ $56.1$ $57.8$ $58.7$ $50.9$ $65.9$ $66.6$ $71.2$ $39.2$ $48.0$ $52.1$ $53.8$
CRB [49] CVPR’22 - $58.7$ $59.1$ $60.9$ $54.2$ $56.5$ $58.9$
LaserMix [26] CVPR’23 $50.6$ $60.0$ $61.9$ $62.3$ $55.3$ $69.9$ $71.8$ $73.2$ $44.2$ $53.7$ $55.1$ $56.8$
LiM3D [29] CVPR’23 - $61.6$ $62.6$ $62.8$ $\mathbf{60.3}$ $\mathbf{60.5}$ $60.9$
ImageTo360 [44] ICCVW’23 $54.5$ $58.6$ $61.4$ $64.2$
FrustumMix Ours $55.7$ $62.5$ $63.0$ $64.9$ $60.0$ $70.0$ $72.6$ $74.1$ $45.6$ $55.7$ $58.2$ $60.8$
Sup.-only - FRNet (Ours) $44.9$ $60.4$ $61.8$ $63.1$ $51.9$ $68.1$ $70.9$ $74.6$ $42.4$ $53.5$ $55.1$ $57.0$
LaserMix [26] CVPR’23 $52.9$ $62.9$ $63.2$ $65.0$ $58.7$ $71.5$ $72.3$ $75.0$ $45.8$ $56.8$ $57.7$ $59.0$
FrustumMix Ours $\mathbf{55.8}$ $\mathbf{64.8}$ $\mathbf{65.2}$ $\mathbf{65.4}$ $\mathbf{61.2}$ $\mathbf{72.2}$ $\mathbf{74.6}$ $\mathbf{75.4}$ $\mathbf{46.6}$ $57.0$ $59.5$ $\mathbf{61.2}$
Range-Interpolation. Previous range-view methods often suffer from large number of empty pixels, which still exist in FRNet. To alleviate this problem, we propose to interpolate some potential points based on the range image. Specifically, LiDAR points will be first projected to range images as previous works do. We then scan all pixels in the range images and find the empty pixels. For each empty pixel, we will judge whether there are projected points in the pixels along its left and right directions. When both the two pixels have projected points, the interpolated points of this pixel can be calculated by averaging the coordinates of two points. During the training procedure, we also need to generate semantic labels for the interpolated points, which will face two cases. If both right and left points have the same semantic labels, we will set the semantic label of the interpolated point to be the same as them. For another case with different semantic labels, we consider the interpolated points as boundaries of the object, and it will be set to an ignored label, which will not be used to calculate loss.
4 Experiments
4.1 Settings
Datasets. We conduct experiments on four popular LiDAR segmentation datasets. SemanticKITTI [5] consists of $22$ sequences with a total of $43,552$ LiDAR scans. Officially, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) for validation, and sequences $11$ to $21$ ($20,351$ scans) for online testing with $19$ semantic classes. nuScenes [13] provides $1,000$ driving scenes captured with much sparser points by 32-beam LiDAR sensors, and 16 classes are evaluated during inference. SemanticPOSS [37] is a more challenging benchmark collected by Peking University with much smaller and sparser point clouds. It consists of $2,988$ LiDAR scenes with a large number of dynamic instances. Officially, it is divided into $6$ parts, and part $2$ is used to evaluate the performance with $13$ categories. ScribbleKITTI [49] shares the same scenes with SemanticKITTI [5] but only provides weak annotations with line scribbles, which means around $8.06\%$ labeled points available during training.
Evaluation Metrics. In line with standard protocols, we utilize Intersection-over-Union (IoU) and Accuracy (Acc) for each class $i$ and compute the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) to access the performance. IoU and Acc can be calculated as:
$\text{IoU}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}~{},~{}~{}\text{Acc}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}~{},$ (9)
where $\text{TP}_{i},\text{TP}_{i},\text{FN}_{i}$ are true-positives, false-positives, and false-negatives for class $i$, respectively. For robustness evaluation, we follow the practice of Robo3D [25] to use Corruption Error (CE) and Resilience Rate (RR) to evaluate the robustness of FRNet. The CE and RR can be calculated as:
$\text{CE}_{i}=\frac{\sum_{l=1}^{3}(1-\text{IoU}_{i,l})}{\sum_{l=1}^{3}(1-\text{IoU}_{i,l}^{\text{base}})}~{},~{}~{}\text{RR}_{i}=\frac{\sum_{l=1}^{3}\text{IoU}_{i,l}}{3\times\text{IoU}_{\text{clean}}}~{},$ (10)
where $\text{IoU}_{i,l}$ is the task-specific IoU; $\text{IoU}_{i,l}^{\text{base}}$ and $\text{IoU}_{\text{clean}}$ are scores of the baseline model and scores on the clean evaluation set.
Implementation Details. We implement FRNet using the MMDetection3D [11] framework and utilize CENet [8] as our backbone to extract frustum features. We set the shape of the 2D representation to $64\times 512$ for SemanticKITTI [5] and ScribbelKITTI [49] and $32\times 480$ for nuScenes [13] and SemanticPOSS [37], respectively. We choose AdamW [34] as our default optimizer and utilize the OneCycle scheduler [45] with $lr=0.01$. The models are trained for $50$ epochs on SemanticKITTI [5] and ScribbelKITTI [49], and $80$ epochs on nuScenes [13] and SemanticPOSS [37], respectively, with a batch size of $2$ with $4$ GPUs.
4.2 Comparative Study
Comparisons to State of the Arts. Firstly, we compare the proposed FRNet with state-of-the-art LiDAR segmentation methods from different representations on both fully-supervised and weakly-supervised benchmarks. Tab. 1 summarizes all the results and the FPS of all the models is evaluated with a single GeForce RTX 2080Ti GPU for fair comparison. Fast-FRNet takes resnet18 [19] as the 2D backbone and the shapes of 2D representation on four datasets are $32\times 360$. Compared with multi-view methods such as SphereFormer [27] and UniSeg [32], although FRNet does not have a significant advantage in performance, its parameters and FPS are far superior to theirs. Compared with voxel-view and range-view methods, FRNet achieves appealing performance while still maintaining high efficiency, and Fast-FRNet achieves the fastest inference speed with only a little drop in performance. Fig. 6 also shows qualitative results among state-of-the-art methods, including CENet [8] and RangeViT [4] on the val set of SemanticKITTI [5]. FRNet reduces prediction errors over large areas, especially for vegetation and parking.
Label-Efficient LiDAR Segmentation. Recently, semi-supervised learning has been explored in LiDAR segmentation. In this work, we follow the paradigm of LaserMix [26] to conduct FRNet on the semi-supervised segmentation task. Here, we use FRNet as the backbone and the proposed FrustumMix to mix labeled and unlabeled. Tab. 2 summarizes the results of recent methods. Results show that our FRNet backbone achieves appealing performance among state-of-the-art methods. Meanwhile, we compare the proposed FrustumMix with LaserMix among different backbones, including Cylinder3D [65] and FRNet. Results show that the proposed FrustumMix outperforms LaserMix, even in voxel representation.
Robustness. Robustness reflects the model’s ability to generalization to a certain extent, which is particularly important in autonomous driving, especially in extreme weather conditions. To this end, we also evaluate the robustness of FRNet using the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [25], which include 8 different conditions and each condition has 3 different levels. The models trained with SemanticKITTI [5] and nuScenes [13] are used to evaluate the performance on SemanticKITTI-C and nuScenes-C [25], respectively. As shown in Tab. 3, FRNet demonstrates strong robustness, especially under extreme weather conditions.
[…]
Method Venue SemKITTI-C nuScenes-C
mCE normal-$\downarrow$ mRR normal-$\uparrow$ mCE normal-$\downarrow$ mRR normal-$\uparrow$
MinkUNet [10] CVPR’19 $100.0$
$\mathbf{81.9}$
$100.0$
$74.4$
Cylinder3D [65] CVPR’21
$103.3$
$80.1$
$111.8$
$72.9$
2DPASS [56] ECCV’22
$106.1$
$77.5$
$\mathbf{98.6}$
$75.2$
FIDNet [64] IROS’20
$113.8$
$77.0$
$122.4$
$73.3$
CENet [8] ICME’22
$103.4$
$81.3$
$112.8$
$76.0$
FRNet Ours
$\mathbf{96.8}$
$80.0$
$\mathbf{98.6}$
$\mathbf{77.5}$
Figure 5: Ablation study on different configurations in FRNet. Tables show the comparisons of: a) Different range-view backbones on the val set of SemanticKITTI [5]. b) Different mixing strategies on the val set of nuScenes [13]. c) Different mixing strategies on the val set of SemanticKITTI [5]. Color-filled boxes denote mAcc while empty boxes denote mIoU in (b) and (c).
Figure 6: Qualitative results among state-of-the-art LiDAR segmentation methods [8, 4]. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
4.3 Ablation Study
In this section, we will discuss each component in FRNet on the val set of SemanticKITTI [5] and nuScenes [13].
Model Analysis. We first analyze different components in FRNet, including the Frustum Feature Encoder, Frustum-Point (FP) Fusion, Fusion Head, Frustum-level Supervision, and Range-Interpolation. The results are summarized in Tab. 4. As the baseline of FRNet, the Frustum Feature Encoder module extracts per-point features in the frustum region and predicts semantic labels of each point with the supervision of point-level labels, and the performance is $62.7\%$ and $72.8\%$ on SemanticKITTI [5] and nuScenes [13], respectively. When we use the FP Fusion module to update per-point features hierarchically, the performance improves by $1.6\%$ and $0.8\%$ on the two datasets, respectively. With the frustum-level supervision, the performance further improves by $1.9\%$ and $2.8\%$. Next, we utilize a Fusion Head module to fuse different level features for predicting, and it improves by $0.8\%$ on both datasets. To alleviate the empty pixels in 2D representations, we propose a Range-Interpolation method to generate some potential points. Results show that this augmentation can also help to improve the performance by $0.6\%$ and $0.5\%$. Finally, following most works, we also use Test Time Augmentation (TTA) during inference, and it will bring an improvement of $1.1\%$ and $1.3\%$, respectively.
[…]
FFE FP FH FS RI TTA SemKITTI nuScenes
mIoU mAcc mIoU mAcc
$62.7$
$72.3$
$72.8$
$82.7$
$64.3$
$73.5$
$73.6$
$83.4$
$66.2$
$73.9$
$76.4$
$84.6$
$67.0$
$74.1$
$77.2$
$84.8$
$67.6$
$74.2$
$77.7$
$\mathbf{85.2}$
$\mathbf{68.7}$
$\mathbf{74.9}$
$\mathbf{79.0}$
$85.0$
Mixing Strategy. Although we have validated the effectiveness of our proposed FrustumMix in semi-supervised segmentation, we still compare FrustumMix with different popular outdoor mix strategies, including LaserMix [26], PolarMix [52]. As shown in Fig. 5 (b) and (c), the proposed FrustumMix significantly improves the performance and outperforms existing mixing strategies on both datasets.
Range-Interpolation Strategy. Range-Interpolation synthesizes possible points to reduce empty pixels in 2D representations. For the empty pixel in the range image, we initially use a $3\times 3$ window to pick its neighboring pixels and average valid pixel information as the interpolated point. Unfortunately, this results in a $0.4\%$ drop in performance. When we simply average the effective pixels, the generated points might fall out of the expected frustum region due to the empty surroundings. Based on this observation, we re-think the interpolation strategy. For an empty pixel, only when there are projection points on both the left and right or above and below, the calculated interpolation point can fall in the expected region. To this end, we propose two interpolation strategies along horizontal and vertical directions and improve the results by $0.4\%$ and $0.6\%$, respectively. It seems that points along horizontal directions share similar attributes as they line in the same laser beams.
Backbones. To validate the versatility of our method, we select different range-view methods as our 2D backbone to extract frustum features, including SalsaNeXt [12], FIDNet [64], and CENet [8]. As shown in Fig. 5 (a), our proposed frustum-range representation can bring an improvement of $5.0\%$ to $6.8\%$ on different range-view methods.
5 Conclusion
In this work, we presented a scalable Frustum-Range network (FRNet) for LiDAR segmentation, which can be seamlessly integrated into any range-view method. FRNet comprises three main components: the Frustum Feature Encoder Module to extract per-point features in low-level, the Frustum-Point Fusion Module to update per-point features hierarchically, and the Fusion Head Module to fuse different level features to predict more accurate labels. We also introduced two efficient augmentations, FurstumMix and Range-Interpolation, to improve the generalization of FRNet. Extensive experiments showed that our FRNet achieves competitive performance while maintaining high efficiency.
Appendix
[…]
Appendix A Additional Implementation Details
In this section, we provide additional details regarding the implementation and reproduction details of the proposed FRNet and Fast-FRNet.
A.1 Datasets
We adopt the following four LiDAR segmentation datasets to validate the effectiveness and superiority of the proposed approach.
  • SemanticKITTI [5] is a large-scale outdoor dataset which is based on KITTI odometry benchmark [14]. It consists of $22$ sequences, collected from different scenes in Karlsruhe, Germany by Velodyne HDL-64E. Commonly, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) is used for validation, and sequences $11$ to $21$ ($20,351$ scans) are used for testing. The original dataset is annotated with $28$ classes and only $19$ semantic classes are used for single-scan evaluation. The vertical field-of-view is $[-25^{\circ},3^{\circ}]$.
  • nuScenes [13] is a multimodal dataset that is widely used in autonomous driving. The LiDAR point clouds are collected around the streets in Boston and Singapore by Velodyne HDL-32E. It comprises $1,000$ driving scenes, in which $700$ scenes ($28,130$ scans) for training, $150$ scenes ($6,019$ scans) for validation, and $150$ scenes ($6,008$ scans) for testing. The dataset is annotated with $32$ classes and only $16$ semantic classes are used for evaluation. The vertical field-of-view is $[-30^{\circ},10^{\circ}]$.
  • ScribbleKITTI [49] is a scribble annotated dataset with line drawings, which shares the same LiDAR point clouds with SemanticKITTI [5]. It contains $189$ million labeled points, around $8.06\%$ of the total training points. This dataset is suitable for conducting weakly-supervised LiDAR segmentation, which is useful for reducing the annotation cost in existing LiDAR segmentation systems.
  • SemanticPOSS [37] is a more challenging dataset collected in Peking University by Pandora. The dataset includes $6$ sequences. Commonly, sequence $02$ ($500$ scans) is used for validation, and others ($2,488$ scans) are used for training. The dataset is annotated with $14$ classes and only $13$ semantic classes are used for evaluation. The vertical field-of-view is $[-16^{\circ},7^{\circ}]$.
A.2 Hyperparameters
For each point with coordinate $(x,y,z)$, its corresponding frustum region $(u,v)$ in the 2D representation with shape of $(H,W)$ is calculated as:
$\binom{u}{v}=\binom{\frac{1}{2}[1-\arctan(y,x)\pi^{-1}]W}{[1-(\arcsin(zd^{-1})+f_{up})f^{-1}]H}~{},$ (11)
where $d=\sqrt{x^{2}+y^{2}+z^{2}}$ represents the depth of a point, $f=|f_{up}|+|f_{down}|$ represents the vertical field-of-view of the sensor, and $f_{up}$ and $f_{down}$ are the pitch angles at the upward and downward directions, respectively. Points with the same $(u,v)$ are grouped into the same frustum region. The hyperparameter configuration of each module in our proposed FRNet framework is summarized as follows.
  • Frustum Feature Encoder. The input of frustum feature encoder for each point is $(x,y,z,i,d,x-x_{m},y-y_{m},z-z_{m})$, where $(x_{m},y_{m},z_{m})$ is the mean coordinate in the corresponding frustum region. The features are fed into four MLPs that map the channel from $8$ to $64$, $128$, $256$, and $256$, respectively, with batch normalization and ReLU function on the first three layers. Then the point features are pooled into frustum features, which are fed into an MLP that maps the channel from $256$ to $16$ with a ReLU function.
  • 2D Backbone. We adopt a variant of ResNet18 and ResNet34 used in CENet [8] as the backbone of Fast-FRNet and FRNet, respectively. The spatial shape of each stage is $[(H,W),(\frac{H}{2},\frac{W}{2}),(\frac{H}{4},\frac{W}{4}),(\frac{H}{8},\frac{W}{8})]$ and the output channel is $[128,128,128,128]$. The four features are interpolated to the shape of $(H,W)$ for feature concatenation, with two convolutional layers to reduce the channel dimension to $[256,128]$ progressively.
  • Frustum-Point Fusion Module. Point features are updated hierarchically at each stage of the 2D backbone. Accordingly, the channel dimension of point features at each stage is $[128,128,128,128]$. We also concatenate point features at each stage and reduce the channel dimension to $[256,128]$ with two MLPs progressively.
  • Optimization. For point-level supervision, we use the cross-entropy loss. For frustum-level supervision, we use cross-entropy loss, Lovasz-Softmax loss [6], and boundary loss [42], and the weights for each loss are $1.0$, $1.5$, and $1.0$, respectively. The frustum-level loss is also used at each stage of the 2D backbone to supervise the middle frustum features. The weights for point-level loss and frustum-level loss are $1.0$ and $1.0$, respectively.
Algorithm 1 FrustumMix, NumPy-style
# lidar_list: the whole dataset.
# h: height of the 2D representation.
# w: width of the 2D representation.
# f_u: upward pitch angles of the field-of-view.
# f_d: downward pitch angles of the field-of-view.
# num_areas: List of split regions along inclination direction. We set [3, 4, 5, 6] here.
# sample: function to sample the lidar points and labels from the whole dataset.
# group: function to calculate frustum regions for points, refer to Eq. (1).
def FrustumMix(xa, ya, xb, yb):
    # group frustum regions
    ua, va = group(xa, h, w, f_u, f_d)
    ub, vb = group(xb, h, w, f_u, f_d)
    out_points = []
    out_labels = []
    if np.random.rand() > 0.5:
        # mix two scans along inclination direction
        n = np.random.choice(num_areas, size=1)[0]
        inc_list = np.linspace(0, h, n+1, dtype=np.int64)
        for i in range(n):
            start = inc_list[i]
            end = inc_list[i+1]
            if i % 2 == 0:
                idx = (va>=start) & (va<end)
                out_points.append(xa[idx])
                out_labels.append(ya[idx])
            else:
                idx = (vb>=start) & (vb<end)
                out_points.append(xb[idx])
                out_labels.append(yb[idx])
    else:
        # mix two scans along azimuth direction
        start = np.random.randint(0, w//2)
        end = start + w//2
        idxa = (ua<start) | (ua>=end)
        idxb = (ub>=start) & (ub<end)
        out_points = [xa[idxa], xb[idxb]]
        out_labels = [ya[idxa], yb[idxb]]
    out_points = np.concatenate(out_points)
    out_labels = np.concatenate(out_labels)
    return out_points, out_labels
# sample two sans
idx_a = np.random.randint(0, len(lidar_list))
idx_b = np.random.randint(0, len(lidar_list))
points_a, labels_a = sample(idx_a)
points_b, labels_b = sample(idx_b)
# mix the two scans
points, labels = FrustumMix(points_a, labels_a, points_b, labels_b)
Algorithm 2 Range-Interpolation, NumPy-style
# h: height of the range image.
# w: width of the range image.
# f_u: upward pitch angles of the field-of-view.
# f_d: downward pitch angles of the field-of-view.
# points: lidar points with shape [m, c].
# labels: semantic labels with shape: [m,].
# ignore_label: ignored label that is not used to calculate loss.
# laser_scan: function to convert LiDAR points to range image.
def RangeInterpolation(p, l):
    # convert to range images
    # r_m means the valid mask for each pixel. 0 is invalid, while 1 is valid.
    r_p, r_l, r_m = laser_scan(p, l, h, w, f_u, f_d)
    out_points = [p]
    out_labels = [l]
    for y in range(h):
        for x in range(w):
            if r_m[y, x]:
                continue
            if (x-1>=0) and (x+1<w):
                if r_m[y, x-1] and r_m[y, x+1]:
                    # both right and left pixels are valid
                    point = (r_p[y, x-1]+r_p[y, x+1]) / 2
                    if r_l[y, x-1] == r_l[y, x+1]:
                        label = r_l[y, x-1]
                    else:
                        label = ignore_label
                    out_points.append(point)
                    out_labels.append(label)
                    r_p[y, x] = point
                    r_l[y, x] = label
                    r_m[y, x] = 1
    out_points = np.concatenate(out_points)
    out_labels = np.concatenate(out_labels)
    return out_points, out_labels
points, labels = RangeInterpolation(points, labels)
A.3 Training Configurations
The common augmentation methods, including random dropping, random flipping, global rotation, global scale, global translation, and instance copy, are used in FRNet. Additionally, we also use the following proposed augmentation methods to further improve the performance.
  • FrustumMix. Based on the calculated frustum region $(u,v)$, the two scans are randomly mixed along inclination and azimuth directions. For inclination direction, the frustum regions are randomly split into several subsets sampled from $[3,4,5,6]$, while for azimuth direction, the frustum regions are split into $2$ subsets.
  • Range-Interpolation. Range-Interpolation aims to synthesize some possible points based on range images, which helps to fill empty pixels in the 2D representations. We first scan all the pixels in the range image. For each empty pixel, the interpolated points are calculated only when its right and left pixels both have projected points.
During the validation and testing phase, only the Range-Interpolation is kept to fill the empty pixels in 2D representations. To make the final results correspond to the original points, we set the first $N$ labels as the predicted labels as the last multiple points are interpolated points. Moreover, we apply the test time augmentation to improve the performance. Specifically, the input points are flipped along different directions, with randomly rotated, scaled, and jittered. The total number of voted point cloud scenes is $4$. For each scene, FRNet generates confidence scores for each class. We add the confidence scores of the $4$ scenes and then predict the final semantic labels.
Appendix B Additional Experimental Results
In this section, we provide the complete experimental results and qualitative visual examples among the four popular LiDAR segmentation benchmarks used in this work.
(a) Speed vs. Accuracy
(b) Speed vs. Robustness
Figure 7: The scalability analysis of existing LiDAR semantic segmentation approaches. Subfigure (a) The inference speed vs. segmentation accuracy (the higher the better). Subfigure (b) The inference speed vs. corruption error (the lower the better).
B.1 Scalability and Robustness
In this section, we provide additional analysis of existing LiDAR segmentation approaches regarding their trade-offs in-between the inference speed (FPS), segmentation accuracy (mIoU), and out-of-training-distribution robustness (mCE). As shown in Fig. 7, current LiDAR segmentation models pursue either the segmentation accuracy or inference speed. Those more accurate LiDAR segmentation models often contain larger parameter sets and are less efficient during the inference stage. This is particularly predominant for the voxel-based (as shown in yellow circulars) and point-voxel fusion (as shown in blue circulars) approaches. The range-view approaches, on the contrary, are faster in terms of inference speed. However, existing range-view models only achieve sub-par performance compared to the voxel-based and point-voxel fusion approaches. It is worth highlighting again that the proposed FRNet and Fast-FRNet provide a better trade-off in-between speed, accuracy, and robustness. Our models achieve competitive segmentation results and superior robustness over the voxel-based and point-voxel fusion counterparts, while still maintaining high efficiency for real-time LiDAR segmentation.
B.2 Per-Class Results
In this section, we supplement the complete (class-wise) segmentation results of the baselines, prior works, and our proposed FRNet and Fast-FRNet.
B.2.1 SemanticKITTI
Tab. 5 shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [5]. We compare FRNet with different LiDAR representations, including point-view, range-view, voxel-view, and multi-view methods. Results demonstrate that FRNet achieves appealing performance among state-of-the-art methods. Compared with SphereFormer [27] and UniSeg [32], although FRNet obtains sub-par performance, it maintains satisfactory efficiency for real-time processing, achieving the balance between efficiency and accuracy. More results about efficiency and accuracy can be found in the main body. It is worth noting that Fast-FRNet achieves the fastest speed of inference with only a little sacrifice of performance.
B.2.2 nuScenes
Tab. 6 and Tab. 7 show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [13], respectively. The results demonstrate the great advantages of Fast-FRNet and FRNet. For the much sparser LiDAR points, the performance of Fast-FRNet is very close to FRNet with fewer parameters. Compared with popular range-view methods, FRNet achieves great improvement in some dynamic instances, such as car, bicycle, motorcycle, etc.
B.2.3 ScribbleKITTI
To prove the great advantages of FRNet among range-view methods, we also reimplement the popular LiDAR segmentation methods on ScribbleKITTI [49], including RangeNet++ [35], SalsaNext [12], FIRNet [64], CENet [8], and RangeViT [4]. All the models are trained with their official settings. As shown in Tab. 8, FRNet achieves state-of-the-art performance. However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
B.2.4 SemanticPOSS
We also re-implement some popular voxel-view methods, including SPVCNN [47], Cylinder3D [65], and MinkNet [10], on SemanticPOSS [37]. All the experiments are conducted based on MMDetection3D framework [11]. Tab. 9 summarizes the class-wise IoU scores among these methods and some other popular range-view methods. Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods. For some tiny objects, such as traffic sign and cone/stone, FRNet achieves superior improvements.
B.3 Qualitative Results
In this section, we provide additional qualitative examples to further validate the effectiveness and superiority of the proposed FRNet framework.
B.3.1 Visual Comparisons
We supplement visual comparison examples of the proposed FRNet with prior works [4] and show the results in Fig. 9. Compared with range-view methods, FRNet can handle better for objects with complex structures, such as vegetation and terrain. In addition, the objects have much clearer boundaries.
B.3.2 Failure Cases
Although FRNet achieves great improvements among popular range-view methods, there are still some limitations that need to be considered. As shown in Fig. 8, FRNet tends to misclassify objects with similar structures to a large extent, i.e., bicyclist and motorcyclist, sidewalk, and other-ground. Spatial-temporal methods will help to alleviate this problem and yield better performance.
B.3.3 Video Demos
In addition to the figures shown in the main body and supplementary material, we also provide some video demos to support the superiority of FRNet. Specifically, we attached three video clips in the supplementary file, named demo1.mp4, demo2.mp4, and demo3.mp4. The first two demos include the error maps of the proposed FRNet and the prior art RangeViT [4]. The third demo contains the error maps and semantic predictions of FRNet. Each video demo shows LiDAR semantic segmentation sequences with hundreds of frames. All three demos validate again the effectiveness and superiority of FRNet in tackling the challenging LiDAR semantic segmentation task. We hope these video demos can provide a more comprehensive qualitative assessment of our proposed approach.
Table 5: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [5] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist road parking sidewalk other-ground building fence vegetation trunk terrain pole traffic-sign PointNet [39] $14.6$ $46.3$ $1.3$ $0.3$ $0.1$ $0.8$ $0.2$ $0.2$ $0.0$ $61.6$ $15.8$ $35.7$ $1.4$ $41.4$ $12.9$ $31.0$ $4.6$ $17.6$ $2.4$ $3.7$ PointNet++ [40] $20.1$ $53.7$ $1.9$ $0.2$ $0.9$ $0.2$ $0.9$ $1.0$ $0.0$ $72.0$ $18.7$ $41.8$ $5.6$ $62.3$ $16.9$ $46.5$ $13.8$ $30.0$ $6.0$ $8.9$ SqSeg [50] $30.8$ $68.3$ $18.1$ $5.1$ $4.1$ $4.8$ $16.5$ $17.3$ $1.2$ $84.9$ $28.4$ $54.7$ $4.6$ $61.5$ $29.2$ $59.6$ $25.5$ $54.7$ $11.2$ $36.3$ SqSegV2 [51] $39.6$ $82.7$ $21.0$ $22.6$ $14.5$ $15.9$ $20.2$ $24.3$ $2.9$ $88.5$ $42.4$ $65.5$ $18.7$ $73.8$ $41.0$ $68.5$ $36.9$ $58.9$ $12.9$ $41.0$ RandLA-Net [21] $50.3$ $94.0$ $19.8$ $21.4$ $42.7$ $38.7$ $47.5$ $48.8$ $4.6$ $90.4$ $56.9$ $67.9$ $15.5$ $81.1$ $49.7$ $78.3$ $60.3$ $59.0$ $44.2$ $38.1$ RangeNet++ [35] $52.2$ $91.4$ $25.7$ $34.4$ $25.7$ $23.0$ $38.3$ $38.8$ $4.8$ $91.8$ $65.0$ $75.2$ $27.8$ $87.4$ $58.6$ $80.5$ $55.1$ $64.6$ $47.9$ $55.9$ PolarNet [61] $54.3$ $93.8$ $40.3$ $30.1$ $22.9$ $28.5$ $43.2$ $40.2$ $5.6$ $90.8$ $61.7$ $74.4$ $21.7$ $90.0$ $61.3$ $84.0$ $65.5$ $67.8$ $51.8$ $57.5$ MPF [2] $55.5$ $93.4$ $30.2$ $38.3$ $26.1$ $28.5$ $48.1$ $46.1$ $18.1$ $90.6$ $62.3$ $74.5$ $30.6$ $88.5$ $59.7$ $83.5$ $59.7$ $69.2$ $49.7$ $58.1$ 3D-MiniNet [3] $55.8$ $90.5$ $42.3$ $42.1$ $28.5$ $29.4$ $47.8$ $44.1$ $14.5$ $91.6$ $64.2$ $74.5$ $25.4$ $89.4$ $60.8$ $82.8$ $60.8$ $66.7$ $48.0$ $56.6$ SqSegV3 [53] $55.9$ $92.5$ $38.7$ $36.5$ $29.6$ $33.0$ $45.6$ $46.2$ $20.1$ $91.7$ $63.4$ $74.8$ $26.4$ $89.0$ $59.4$ $82.0$ $58.7$ $65.4$ $49.6$ $58.9$ KPConv [48] $58.8$ $96.0$ $32.0$ $42.5$ $33.4$ $44.3$ $61.5$ $61.6$ $11.8$ $88.8$ $61.3$ $72.7$ $31.6$ $95.0$ $64.2$ $84.8$ $69.2$ $69.1$ $56.4$ $47.4$ SalsaNext [12] $59.5$ $91.9$ $48.3$ $38.6$ $38.9$ $31.9$ $60.2$ $59.0$ $19.4$ $91.7$ $63.7$ $75.8$ $29.1$ $90.2$ $64.2$ $81.8$ $63.6$ $66.5$ $54.3$ $62.1$ FIDNet [64] $59.5$ $93.9$ $54.7$ $48.9$ $27.6$ $23.9$ $62.3$ $59.8$ $23.7$ $90.6$ $59.1$ $75.8$ $26.7$ $88.9$ $60.5$ $84.5$ $64.4$ $69.0$ $53.3$ $62.8$ FusionNet [59] $61.3$ $95.3$ $47.5$ $37.7$ $41.8$ $34.5$ $59.5$ $56.8$ $11.9$ $91.8$ $68.8$ $77.1$ $30.8$ $92.5$ $69.4$ $84.5$ $69.8$ $68.5$ $60.4$ $66.5$ PCSCNet [38] $62.7$ $95.7$ $48.8$ $46.2$ $36.4$ $40.6$ $55.5$ $68.4$ $55.9$ $89.1$ $60.2$ $72.4$ $23.7$ $89.3$ $64.3$ $84.2$ $68.2$ $68.1$ $60.5$ $63.9$ KPRNet [23] $63.1$ $95.5$ $54.1$ $47.9$ $23.6$ $42.6$ $65.9$ $65.0$ $16.5$ $93.2$ $73.9$ $80.6$ $30.2$ $91.7$ $68.4$ $85.7$ $69.8$ $71.2$ $58.7$ $64.1$ TornadoNet [16] $63.1$ $94.2$ $55.7$ $48.1$ $40.0$ $38.2$ $63.6$ $60.1$ $34.9$ $89.7$ $66.3$ $74.5$ $28.7$ $91.3$ $65.6$ $85.6$ $67.0$ $71.5$ $58.0$ $65.9$ LiteHDSeg [43] $63.8$ $92.3$ $40.0$ $55.4$ $37.7$ $39.6$ $59.2$ $71.6$ $54.3$ $93.0$ $68.2$ $78.3$ $29.3$ $91.5$ $65.0$ $78.2$ $65.8$ $65.1$ $59.5$ $67.7$ RangeViT [4] $64.0$ $95.4$ $55.8$ $43.5$ $29.8$ $42.1$ $63.9$ $58.2$ $38.1$ $93.1$ $70.2$ $80.0$ $32.5$ $92.0$ $69.0$ $85.3$ $70.6$ $71.2$ $60.8$ $64.7$ CENet [8] $64.7$ $91.9$ $58.6$ $50.3$ $40.6$ $42.3$ $68.9$ $65.9$ $43.5$ $90.3$ $60.9$ $75.1$ $31.5$ $91.0$ $66.2$ $84.5$ $69.7$ $70.0$ $61.5$ $67.6$ SVASeg [63] $65.2$ $96.7$ $56.4$ $57.0$ $49.1$ $56.3$ $70.6$ $67.0$ $15.4$ $92.3$ $65.9$ $76.5$ $23.6$ $91.4$ $66.1$ $85.2$ $72.9$ $67.8$ $63.9$ $65.2$ AMVNet [31] $65.3$ $96.2$ $59.9$ $54.2$ $48.8$ $45.7$ $71.0$ $65.7$ $11.0$ $90.1$ $71.0$ $75.8$ $32.4$ $92.4$ $69.1$ $85.6$ $71.7$ $69.6$ $62.7$ $67.2$ GFNet [41] $65.4$ $96.0$ $53.2$ $48.3$ $31.7$ $47.3$ $62.8$ $57.3$ $44.7$ $\mathbf{93.6}$ $72.5$ $\mathbf{80.8}$ $31.2$ $\mathbf{94.0}$ $\mathbf{73.9}$ $85.2$ $71.1$ $69.3$ $61.8$ $68.0$ JS3C-Net [55] $66.0$ $95.8$ $59.3$ $52.9$ $54.3$ $46.0$ $69.5$ $65.4$ $39.9$ $88.9$ $61.9$ $72.1$ $31.9$ $92.5$ $70.8$ $84.5$ $69.8$ $67.9$ $60.7$ $68.7$ MaskRange [17] $66.1$ $94.2$ $56.0$ $55.7$ $59.2$ $52.4$ $67.6$ $64.8$ $31.8$ $91.7$ $70.7$ $77.1$ $29.5$ $90.6$ $65.2$ $84.6$ $68.5$ $69.2$ $60.2$ $66.6$ SPVNAS [47] $66.4$ $97.3$ $51.5$ $50.8$ $59.8$ $58.8$ $65.7$ $65.2$ $43.7$ $90.2$ $67.6$ $75.2$ $16.9$ $91.3$ $65.9$ $86.1$ $73.4$ $71.0$ $64.2$ $66.9$ MSSNet [46] $66.7$ $96.8$ $52.2$ $48.5$ $54.4$ $56.3$ $67.0$ $70.9$ $49.3$ $90.1$ $65.5$ $74.9$ $30.2$ $90.5$ $64.9$ $84.9$ $72.7$ $69.2$ $63.2$ $65.1$ Cylinder3D [65] $68.9$ $97.1$ $67.6$ $63.8$ $50.8$ $58.5$ $73.7$ $69.2$ $48.0$ $92.2$ $65.0$ $77.0$ $32.3$ $90.7$ $66.5$ $85.6$ $72.5$ $69.8$ $62.4$ $66.2$ AF2S3Net [9] $69.7$ $94.5$ $65.4$ $\mathbf{86.8}$ $39.2$ $41.1$ $\mathbf{80.7}$ $80.4$ $74.3$ $91.3$ $68.8$ $72.5$ $\mathbf{53.5}$ $87.9$ $63.2$ $70.2$ $68.5$ $53.7$ $61.5$ $71.0$ RPVNet [54] $70.3$ $97.6$ $68.4$ $68.7$ $44.2$ $61.1$ $75.9$ $74.4$ $73.4$ $93.4$ $70.3$ $80.7$ $33.3$ $93.5$ $72.1$ $86.5$ $75.1$ $71.7$ $64.8$ $61.4$ SDSeg3D [28] $70.4$ $97.4$ $58.7$ $54.2$ $54.9$ $65.2$ $70.2$ $74.4$ $52.2$ $90.9$ $69.4$ $76.7$ $41.9$ $93.2$ $71.1$ $86.1$ $74.3$ $71.1$ $65.4$ $70.6$ GASN [58] $70.7$ $96.9$ $65.8$ $58.0$ $59.3$ $61.0$ $80.4$ $\mathbf{82.7}$ $46.3$ $89.8$ $66.2$ $74.6$ $30.1$ $92.3$ $69.6$ $87.3$ $73.0$ $72.5$ $66.1$ $71.6$ PVKD [20] $71.2$ $97.0$ $67.9$ $69.3$ $53.5$ $60.2$ $75.1$ $73.5$ $50.5$ $91.8$ $70.9$ $77.5$ $41.0$ $92.4$ $69.4$ $86.5$ $73.8$ $71.9$ $64.9$ $65.8$ Fast-FRNet $72.5$ $97.1$ $66.0$ $73.1$ $59.3$ $65.7$ $76.0$ $78.9$ $54.5$ $91.8$ $72.6$ $77.6$ $42.0$ $92.4$ $70.6$ $86.5$ $71.9$ $72.5$ $63.1$ $66.7$ 2DPASS [56] $72.9$ $97.0$ $63.6$ $63.4$ $61.1$ $61.5$ $77.9$ $81.3$ $74.1$ $89.7$ $67.4$ $74.7$ $40.0$ $93.5$ $72.9$ $86.2$ $73.9$ $71.0$ $65.0$ $70.4$ RangeFormer [24] $73.3$ $96.7$ $69.4$ $73.7$ $59.9$ $66.2$ $78.1$ $75.9$ $58.1$ $92.4$ $73.0$ $78.8$ $42.4$ $92.3$ $70.1$ $86.6$ $73.3$ $72.8$ $66.4$ $66.6$ FRNet $73.3$ $97.3$ $67.9$ $74.6$ $59.4$ $66.3$ $78.1$ $79.2$ $57.3$ $92.1$ $73.0$ $78.1$ $41.8$ $92.7$ $71.0$ $86.7$ $73.2$ $72.5$ $64.7$ $67.3$ SphereFormer [27] $74.8$ $97.5$ $70.1$ $70.5$ $59.6$ $67.7$ $79.0$ $80.4$ $\mathbf{75.3}$ $91.8$ $69.7$ $78.2$ $41.3$ $93.8$ $72.8$ $86.7$ $75.1$ $72.4$ $66.8$ $\mathbf{72.9}$ UniSeg [32] $\mathbf{75.2}$ $\mathbf{97.9}$ $\mathbf{71.9}$ $75.2$ $\mathbf{63.6}$ $\mathbf{74.1}$ $78.9$ $74.8$ $60.6$ $92.6$ $\mathbf{74.0}$ $79.5$ $46.1$ $93.4$ $72.7$ $\mathbf{87.5}$ $\mathbf{76.3}$ $\mathbf{73.1}$ $\mathbf{68.3}$ $68.5$
Table 6: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [13]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation AF2S3Net [9] $62.2$ $60.3$ $12.6$ $82.3$ $80.0$ $20.1$ $62.0$ $59.0$ $49.0$ $42.2$ $67.4$ $94.2$ $68.0$ $64.1$ $68.6$ $82.9$ $82.4$ RangeNet++ [35] $65.5$ $66.0$ $21.3$ $77.2$ $80.9$ $30.2$ $66.8$ $69.6$ $52.1$ $54.2$ $72.3$ $94.1$ $66.6$ $63.5$ $70.1$ $83.1$ $79.8$ PolarNet [61] $71.0$ $74.7$ $28.2$ $85.3$ $90.9$ $35.1$ $77.5$ $71.3$ $58.8$ $57.4$ $76.1$ $96.5$ $71.1$ $74.7$ $74.0$ $87.3$ $85.7$ PCSCNet [38] $72.0$ $73.3$ $42.2$ $87.8$ $86.1$ $44.9$ $82.2$ $76.1$ $62.9$ $49.3$ $77.3$ $95.2$ $66.9$ $69.5$ $72.3$ $83.7$ $82.5$ SalsaNext [12] $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $63.1$ $61.3$ $76.5$ $96.0$ $70.8$ $71.2$ $71.5$ $86.7$ $84.4$ SVASeg [63] $74.7$ $73.1$ $44.5$ $88.4$ $86.6$ $48.2$ $80.5$ $77.7$ $65.6$ $57.5$ $82.1$ $96.5$ $70.5$ $74.7$ $74.6$ $87.3$ $86.9$ RangeViT [4] $75.2$ $75.5$ $40.7$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $66.3$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $89.9$ $87.2$ Cylinder3D [65] $76.1$ $76.4$ $40.3$ $91.2$ $\mathbf{93.8}$ $51.3$ $78.0$ $78.9$ $64.9$ $62.1$ $84.4$ $96.8$ $71.6$ $76.4$ $75.4$ $90.5$ $87.4$ AMVNet [31] $76.1$ $\mathbf{79.8}$ $32.4$ $82.2$ $86.4$ $\mathbf{62.5}$ $81.9$ $75.3$ $\mathbf{72.3}$ $\mathbf{83.5}$ $65.1$ $\mathbf{97.4}$ $67.0$ $\mathbf{78.8}$ $74.6$ $90.8$ $87.9$ RPVNet [54] $77.6$ $78.2$ $43.4$ $92.7$ $93.2$ $49.0$ $85.7$ $80.5$ $66.0$ $66.9$ $84.0$ $96.9$ $73.5$ $75.9$ $70.6$ $90.6$ $88.9$ RangeFormer [24] $78.1$ $78.0$ $\mathbf{45.2}$ $94.0$ $92.9$ $58.7$ $83.9$ $77.9$ $69.1$ $63.7$ $85.6$ $96.7$ $74.5$ $75.1$ $75.3$ $89.1$ $87.5$ SphereFormer [27] $78.4$ $77.7$ $43.8$ $94.5$ $93.1$ $52.4$ $\mathbf{86.9}$ $\mathbf{81.2}$ $65.4$ $73.4$ $85.3$ $97.0$ $73.4$ $75.4$ $75.0$ $\mathbf{91.0}$ $\mathbf{89.2}$ Fast-FRNet $78.8$ $78.7$ $42.3$ $\mathbf{95.6}$ $93.1$ $58.9$ $86.3$ $77.9$ $66.9$ $72.1$ $85.4$ $97.0$ $\mathbf{76.3}$ $76.5$ $76.2$ $89.7$ $87.8$ FRNet $\mathbf{79.0}$ $78.5$ $43.9$ $95.4$ $93.2$ $56.3$ $85.8$ $79.0$ $68.5$ $72.8$ $\mathbf{86.5}$ $97.1$ $75.9$ $77.0$ $\mathbf{76.4}$ $89.7$ $88.0$
Table 7: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [13]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation PolarNet [61] $69.4$ $72.2$ $16.8$ $77.0$ $86.5$ $51.1$ $69.7$ $64.8$ $54.1$ $69.7$ $63.5$ $96.6$ $67.1$ $77.7$ $72.1$ $87.1$ $84.5$ JS3C-Net [55] $73.6$ $80.1$ $26.2$ $87.8$ $84.5$ $55.2$ $72.6$ $71.3$ $66.3$ $76.8$ $71.2$ $96.8$ $64.5$ $76.9$ $74.1$ $87.5$ $86.1$ PMF [66] $77.0$ $82.0$ $40.0$ $81.0$ $88.0$ $64.0$ $79.0$ $80.0$ $76.0$ $81.0$ $67.0$ $97.0$ $68.0$ $78.0$ $74.0$ $90.0$ $88.0$ Cylinder3D [65] $77.2$ $82.8$ $29.8$ $84.3$ $89.4$ $63.0$ $79.3$ $77.2$ $73.4$ $84.6$ $69.1$ $97.7$ $70.2$ $80.3$ $75.5$ $90.4$ $87.6$ AMVNet [31] $77.3$ $80.6$ $32.0$ $81.7$ $88.9$ $67.1$ $84.3$ $76.1$ $73.5$ $84.9$ $67.3$ $97.5$ $67.4$ $79.4$ $75.5$ $91.5$ $88.7$ SPVCNN [47] $77.4$ $80.0$ $30.0$ $91.9$ $90.8$ $64.7$ $79.0$ $75.6$ $70.9$ $81.0$ $74.6$ $97.4$ $69.2$ $80.0$ $76.1$ $89.3$ $87.1$ AF2S3Net [9] $78.3$ $78.9$ $52.2$ $89.9$ $84.2$ $77.4$ $74.3$ $77.3$ $72.0$ $83.9$ $73.8$ $97.1$ $66.5$ $77.5$ $74.0$ $87.7$ $86.8$ 2D3DNet [15] $80.0$ $83.0$ $59.4$ $88.0$ $85.1$ $63.7$ $84.4$ $82.0$ $76.0$ $84.8$ $71.9$ $96.9$ $67.4$ $79.8$ $76.0$ $92.1$ $89.2$ RangeFormer [24] $80.1$ $85.6$ $47.4$ $91.2$ $90.9$ $70.7$ $84.7$ $77.1$ $74.1$ $83.2$ $72.6$ $97.5$ $70.7$ $79.2$ $75.4$ $91.3$ $88.9$ GASN [58] $80.4$ $85.5$ $43.2$ $90.5$ $92.1$ $64.7$ $86.0$ $83.0$ $73.3$ $83.9$ $75.8$ $97.0$ $71.0$ $81.0$ $\mathbf{77.7}$ $91.6$ $\mathbf{90.2}$ 2DPASS [56] $80.8$ $81.7$ $55.3$ $92.0$ $91.8$ $73.3$ $86.5$ $78.5$ $72.5$ $84.7$ $75.5$ $97.6$ $69.1$ $79.9$ $75.5$ $90.2$ $88.0$ LidarMultiNet [57] $81.4$ $80.4$ $48.4$ $94.3$ $90.0$ $71.5$ $87.2$ $\mathbf{85.2}$ $\mathbf{80.4}$ $86.9$ $74.8$ $97.8$ $67.3$ $80.7$ $76.5$ $92.1$ $89.6$ SphereFormer [27] $81.9$ $83.3$ $39.2$ $\mathbf{94.7}$ $\mathbf{92.5}$ $77.5$ $84.2$ $84.4$ $79.1$ $\mathbf{88.4}$ $\mathbf{78.3}$ $\mathbf{97.9}$ $69.0$ $\mathbf{81.5}$ $77.2$ $\mathbf{93.4}$ $\mathbf{90.2}$ Fast-FRNet $82.1$ $85.3$ $62.9$ $92.1$ $91.5$ $76.5$ $87.7$ $75.9$ $73.2$ $86.2$ $76.0$ $97.8$ $\mathbf{71.9}$ $80.9$ $77.1$ $90.8$ $88.1$ FRNet $82.5$ $85.8$ $65.4$ $92.1$ $91.6$ $77.4$ $87.9$ $77.4$ $74.3$ $86.0$ $75.7$ $97.8$ $71.8$ $80.8$ $77.0$ $91.0$ $88.3$ UniSeg [32] $\mathbf{83.5}$ $\mathbf{85.9}$ $\mathbf{71.2}$ $92.1$ $91.6$ $\mathbf{80.5}$ $\mathbf{88.0}$ $80.9$ $76.0$ $86.3$ $76.7$ $97.7$ $71.8$ $80.7$ $76.7$ $91.3$ $88.8$
Table 8: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [49] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist road parking sidewalk other-ground building fence vegetation trunk terrain pole traffic-sign RangeNet++ [35] $44.6$ $84.6$ $24.6$ $38.9$ $13.7$ $12.9$ $29.0$ $51.4$ $0.0$ $86.0$ $35.2$ $72.4$ $4.5$ $80.2$ $41.6$ $80.1$ $51.2$ $66.9$ $43.9$ $34.5$ SalsaNext [12] $50.3$ $85.6$ $33.8$ $42.7$ $8.4$ $20.8$ $64.3$ $71.5$ $0.2$ $85.6$ $31.6$ $72.0$ $1.1$ $84.0$ $39.9$ $81.0$ $62.0$ $63.6$ $60.5$ $46.3$ RangeViT [4] $53.6$ $85.6$ $31.6$ $50.1$ $40.3$ $36.3$ $57.6$ $68.7$ $0.0$ $86.1$ $32.6$ $75.2$ $0.3$ $87.9$ $49.3$ $83.6$ $62.8$ $67.5$ $59.6$ $43.7$ FIDNet [64] $54.1$ $85.6$ $36.7$ $48.7$ $60.8$ $38.4$ $63.3$ $68.2$ $0.0$ $84.1$ $25.9$ $71.2$ $0.4$ $85.6$ $41.3$ $81.7$ $64.1$ $62.7$ $61.5$ $48.0$ MinkNet [10] $55.0$ $88.1$ $13.2$ $55.1$ $72.3$ $36.9$ $61.3$ $77.1$ $0.0$ $83.4$ $32.7$ $71.0$ $0.3$ $90.0$ $50.0$ $84.1$ $66.6$ $65.8$ $61.6$ $35.2$ CENet [8] $55.7$ $86.1$ $39.4$ $53.2$ $61.0$ $46.1$ $69.2$ $72.2$ $0.0$ $85.7$ $28.7$ $72.6$ $1.1$ $85.8$ $43.1$ $81.8$ $64.2$ $63.8$ $59.6$ $45.0$ SPVCNN [47] $56.9$ $88.6$ $25.7$ $55.9$ $67.4$ $48.8$ $65.0$ $78.2$ $0.0$ $82.6$ $30.4$ $70.1$ $0.3$ $\mathbf{90.5}$ $49.6$ $84.4$ $\mathbf{67.6}$ $66.1$ $61.6$ $\mathbf{48.7}$ Cylinder3D [65] $57.0$ $88.5$ $39.9$ $58.0$ $58.4$ $48.1$ $68.6$ $77.0$ $\mathbf{0.5}$ $84.4$ $30.4$ $72.2$ $2.5$ $89.4$ $48.4$ $81.9$ $64.6$ $59.8$ $61.2$ $\mathbf{48.7}$ Fast-FRNet $62.4$ $90.8$ $41.0$ $\mathbf{66.8}$ $81.7$ $\mathbf{64.0}$ $70.5$ $84.2$ $0.0$ $91.3$ $35.1$ $78.3$ $0.0$ $89.4$ $64.4$ $85.0$ $65.5$ $69.5$ $60.5$ $47.7$ RangeFormer [24] $63.0$ $\mathbf{92.6}$ $\mathbf{51.6}$ $65.7$ $74.4$ $49.6$ $71.6$ $82.1$ $0.0$ $\mathbf{94.8}$ $\mathbf{44.4}$ $\mathbf{80.6}$ $\mathbf{11.4}$ $85.6$ $56.9$ $\mathbf{87.2}$ $64.1$ $\mathbf{77.0}$ $\mathbf{62.7}$ $45.1$ FRNet $\mathbf{63.1}$ $90.5$ $42.3$ $\mathbf{66.8}$ $\mathbf{82.4}$ $63.0$ $\mathbf{73.4}$ $\mathbf{86.2}$ $0.0$ $92.1$ $35.1$ $79.1$ $0.3$ $90.4$ $\mathbf{64.7}$ $85.4$ $65.8$ $70.7$ $61.8$ $48.6$
Appendix C Broader Impact
In this section, we elaborate on the positive societal influence and potential limitations of the proposed FRNet.
C.1 Positive Societal Impacts
Conducting efficient and accurate range view LiDAR semantic segmentation for autonomous driving has several positive societal impacts, particularly in terms of enhancing road safety, improving transportation efficiency, and contributing to broader technological advancements.
  • Enhanced Road Safety. One of the most significant benefits is the improvement in road safety. Autonomous vehicles equipped with advanced LiDAR semantic segmentation can accurately detect and classify objects in their environment, such as other vehicles, pedestrians, cyclists, and road obstacles. This precise perception capability allows for safer navigation and decision-making, potentially reducing accidents caused by human error.
  • Reduced Traffic Congestion. Autonomous vehicles with advanced perception systems can optimize driving patterns, leading to smoother traffic flow. This can significantly reduce traffic congestion, especially in urban areas, thereby saving time for commuters and reducing stress associated with driving.
  • Emergency Response and Healthcare. In emergency situations, autonomous vehicles can be used to quickly and safely transport patients or deliver medical supplies. The precision of LiDAR segmentation ensures that these vehicles can navigate through complex environments effectively.
  • Advancements in Smart City Infrastructure. The integration of autonomous vehicles with smart city initiatives can lead to more efficient urban planning and infrastructure development. LiDAR data can be used not just for navigation but also for gathering urban data, which can inform city planning and maintenance.
C.2 Potential Limitations
Although the proposed FRNet and Fast-FRNet are capable of providing a better trade-off between LiDAR segmentation network efficiency and accuracy, there are several limitations within the current framework. First, frustum-level supervision counts the high-frequency semantic labels in the frustum region, which will cover the objects with few points. Such supervision weakens the performance of FRNet on tiny objects. Second, FRNet cannot handle objects with similar structures well. Some objects share a similar appearance with different semantic attributes, which limits the ability of FRNet to distinguish them in a 2D manner.
Appendix D Public Resources Used
In this section, we acknowledge the use of public resources, during the course of this work.
D.1 Public Codebase Used
We acknowledge the use of the following public codebase, during the course of this work:
D.2 Public Datasets Used
We acknowledge the use of the following public datasets, during the course of this work:
D.3 Public Implementations Used
We acknowledge the use of the following public implementations, during the course of this work:
Table 9: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [37] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method (year)
mIoU
person
rider
car
truck
plants
traffic sign
pole
trashcan
building
cone/stone
walk
fence
bike
SqSeg [50] [’18] $18.9$ $14.2$ $1.0$ $13.2$ $10.4$ $28.0$ $5.1$ $5.7$ $2.3$ $43.6$ $0.2$ $15.6$ $31.0$ $75.0$
SqSegV2 [51] [’19] $30.0$ $48.0$ $9.4$ $48.5$ $11.3$ $50.1$ $6.7$ $6.2$ $14.8$ $60.4$ $5.2$ $22.1$ $36.1$ $71.3$
RangeNet++ [35] [’19] $30.9$ $57.3$ $4.6$ $35.0$ $14.1$ $58.3$ $3.9$ $6.9$ $24.1$ $66.1$ $6.6$ $23.4$ $28.6$ $73.5$
MINet [30] [’21] $43.2$ $62.4$ $12.1$ $63.8$ $22.3$ $68.6$ $16.7$ $30.1$ $28.9$ $75.1$ $28.6$ $32.2$ $44.9$ $76.3$
FIDNet [64] [’21] $46.4$ $72.2$ $23.1$ $72.7$ $23.0$ $68.0$ $22.2$ $28.6$ $16.3$ $73.1$ $34.0$ $40.9$ $50.3$ $79.1$
SPVCNN [47] [’20] $48.4$ $72.5$ $24.7$ $72.1$ $31.4$ $72.7$ $10.8$ $41.3$ $31.8$ $78.4$ $23.8$ $42.6$ $51.7$ $75.3$
CENet [8] [’22] $50.3$ $75.5$ $22.0$ $77.6$ $25.3$ $72.2$ $18.2$ $31.5$ $\mathbf{48.1}$ $76.3$ $27.7$ $47.7$ $51.4$ $\mathbf{80.3}$
Fast-FRNet $52.5$ $76.9$ $28.3$ $79.9$ $28.8$ $73.8$ $30.2$ $32.9$ $28.5$ $80.9$ $40.7$ $47.3$ $55.9$ $79.0$
Cylinder3D [65] [’21] $52.9$ $75.9$ $\mathbf{30.0}$ $75.8$ $28.7$ $\mathbf{75.7}$ $29.5$ $37.2$ $36.7$ $\mathbf{82.3}$ $34.1$ $47.5$ $53.9$ $80.1$
MinkNet [10] [’19] $53.1$ $\mathbf{77.8}$ $29.1$ $\mathbf{81.3}$ $\mathbf{33.9}$ $75.2$ $22.0$ $\mathbf{42.5}$ $36.4$ $80.7$ $23.9$ $\mathbf{51.2}$ $\mathbf{57.5}$ $79.1$
FRNet $\mathbf{53.5}$ $\underline{77.7}$ $28.7$ $81.2$ $28.9$ $74.3$ $\mathbf{30.4}$ $34.5$ $29.8$ $81.1$ $\mathbf{45.6}$ $47.9$ $56.4$ $79.2$
Figure 8: Failure case examples in FRNet. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
Figure 9: Qualitative comparisons with state-of-the-art LiDAR segmentation methods [4] on the val set of nuScenes [13]. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.